A speed boat can travel at a rate of 40 mph. At this rate, what is the distance that the speedboat that the speedboat will travel in 6 minutes
Answer by jim_thompson5910(35256)

6 min = 6/60 = 1/10 hr


d = rt

d = 40(1/10)

d = 40/10

d = 4


So the boat will travel 4 miles.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim